OpenStudy (anonymous):

Yesterday, Carolyn's bank account balance was –$31.16. She deposited some money into her account so that the balance is now $15.82. How much did she deposit into her account? A. $15.34 B. $16.34 C. $46.98 D. $47.02

OpenStudy (phi):

if you call the unknown amount x you have –$31.16 + x = $15.82

OpenStudy (phi):

to solve for x, add 31.16 to both sides -31.16 + 31.16 + x = 15.82+31.16 and simplify

OpenStudy (anonymous):

so whats the answer????

OpenStudy (khaakenajaf):

A

OpenStudy (anonymous):

Thanks!

OpenStudy (khaakenajaf):

no prob

OpenStudy (anonymous):

c